OpenStudy (anonymous):

Michele Solve 4 log12 2 + log12 x = log12 96.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

x=12 yes

OpenStudy (anonymous):

michele

OpenStudy (michele_laino):

here we can rewrite the left side as below: \[\Large \begin{gathered} 4{\log _{12}}2 + {\log _{12}}x = {\log _{12}}{2^4} + {\log _{12}}x = \hfill \\ = {\log _{12}}16 + {\log _{12}}x = {\log _{12}}\left( {16x} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

yes x=12 good

OpenStudy (anonymous):

thank michele

OpenStudy (campbell_st):

use the log laws for powers here is an example \[alog(b) = \log(b^a)\] this is the 1st thing you need to do with the 1st term... then simplify it

OpenStudy (anonymous):

Solve 2 log2 2 + 2 log2 6 –log2 3x = 3 michele this

OpenStudy (michele_laino):

please wait

OpenStudy (michele_laino):

your equation becomes: \[{\log _{12}}\left( {16x} \right) = {\log _{12}}1296\]

OpenStudy (campbell_st):

then use the log law for multiplication \[\log(a) + \log(b) = \log(ab)\] that needs to be used for both terms on the left hand side

OpenStudy (campbell_st):

then you will have 2 log terms with the same base that can be equated and solved

OpenStudy (anonymous):

can i be right x=12

OpenStudy (michele_laino):

no, I think that is not the right answer

OpenStudy (campbell_st):

no its not 12.... read the information and equate the expressions 16x = 96 solve for x

OpenStudy (anonymous):

96 is right michele?

OpenStudy (michele_laino):

using the uniqueness of logarithm we have to solve this equation: \[16x = 96\]

OpenStudy (michele_laino):

so what is: 96/16=...?

OpenStudy (anonymous):

fraction this time again!

OpenStudy (anonymous):

6

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

good

OpenStudy (anonymous):

one more michele thank

OpenStudy (michele_laino):

thanks!

OpenStudy (michele_laino):

ok!